40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 554     Entry time: Thu Jul 26 15:33:00 2018
Author: Alena 
Type: General 
Category: General 
Subject: New laser instaled for the old chanber cr0 

Installed a new HeNe laser for the old chamber cr0. No major alignement was required. The laser mount restored the alinement. Will run a test measurement now.

Attachment 1: 20180726_152508.jpg  3.368 MB  Uploaded Thu Jul 26 18:34:40 2018  | Hide | Hide all | Show all
20180726_152508.jpg
Attachment 2: 20180726_152521.jpg  3.853 MB  Uploaded Thu Jul 26 18:35:00 2018  | Hide | Hide all | Show all
20180726_152521.jpg
Attachment 3: 20180726_152530.jpg  2.863 MB  Uploaded Thu Jul 26 18:35:17 2018  | Hide | Hide all | Show all
20180726_152530.jpg
Attachment 4: 20180726_154138.jpg  3.056 MB  Uploaded Thu Jul 26 18:35:29 2018  | Show | Hide all | Show all
ELOG V3.1.3-